# Characterizing and Understanding HGNN Training on GPUs

### v3

@@ 5.2.2. Bandwidth Competition During Data Loading and Synchronization
Fig. 14(a) illustrates the trend of PCIe bus communication bandwidth between the host CPU and GPU devices during distributed training as the number of devices increases. Based on the geometric mean across various model and dataset scenarios, an increase in the number of devices from one to two results in an 8.37% reduction in the average PCIe bus bandwidth connecting the host and devices. This reduction further extends to 22.62% when the number of devices increases to four. The decline in PCIe bandwidth suggests that data transfers between the CPU and GPU, particularly the data loading process, encounter increased latency, consequently diminishing the overall performance of distributed training.
@@
Figure 14. Bandwidth competition between devices: (a) Host-Device PCIe bandwidth; (b) Device-Device NVLink bandwidth.
@@
Fig. 14(b) shows the changes in data transfer bandwidth through NVLink between GPUs in distributed training as the number of devices increases. From the geometric mean across different model and dataset scenarios, when the number of devices increases from one to two, the average NVLink bus bandwidth connecting Device-Device decreases by 26.07%, and when it increases to four devices, the average NVLink data transfer bandwidth decreases by 46.03%. This phenomenon occurs because the NVIDIA Collective Communications Library library, utilized for gradient synchronization in this experiment, employs the All-Reduce algorithm to synchronize gradient information across devices. This process is executed concurrently on multiple devices. As the number of devices increases, the communication network’s complexity escalates, resulting in heightened contention for communication bandwidth among the devices.
@@ 6.1.1. Unique Process of Metapath Instance Generation
Given that HGNNs primarily operate on semantic graphs during execution, it is imperative to construct these semantic graphs from the original HetG, which is unique to HGNNs. Our experimental results for the HAN model on different datasets indicate that the execution time of the SGB stage is, on average, 11.46$\times$ the execution time of one single training epoch. Furthermore, as the length and number of metapaths increase, the execution time of the SGB stage exhibits a marked increase, irrespective of the dataset’s scale. For large-scale HetG datasets, SGB will be a considerably time-consuming stage during in the training process.
@@ 6.1.3. Intricate Two-level Aggregation
GNNs perform a single aggregation step for neighboring vertices within a singular type of relation. In contrast, HGNNs aggregate features from neighbors in each semantic graph generated according to corresponding semantics (relations or metapaths), and then fuse intermediate results of each semantic graph for each vertex. As depicted in Fig. 3, the incorporation of a secondary aggregation level, denoted as the SF stage, incurs an approximate 15% increase in time overhead per epoch.
@@ 6.2.1. More Intricate Hybrid Execution Pattern
Prior work (Yan et al., 2020) highlights the presence of hybrid execution pattern in the execution of the typical GCN model. Specifically, the FP stage (Combination stage in work (Yan et al., 2020)) exhibits a more regular pattern and demonstrates compute-bound execution modes, whereas the NA (Aggregation stage) stage involves numerous random accesses, resulting in memory-bound behavior. According to the analysis in Section 4.1.2, compared to GNNs, HGNNs exhibit more pronounced and complex hybrid execution pattern, which stems from their more intricate model structure.
@@ 7. Optimization Guidelines
In this section, leveraging the findings previously outlined, we provide guidences for optimizing HGNN training from both software and hardware perspectives.
7.1. Software Perspective
@@ 7.1.1. Reasonable Overlapping of Phases
On one hand, a bound-aware kernel fusion method can be proposed to facilitate the overlapping execution of stage with differing bounds. Observations 5 and 6 underscore the presence of intricate hybrid execution patterns during HGNN training. These stages, distinguished by diverse execution bounds, frequently alternate, facilitating overlap execution to harness multiple hardware resources concurrently. This approach enhances overall execution performance by optimizing hardware resource utilization. For example, Graphite (Gong et al., 2022) adopts similar approach to accelerate GNNs on the CPU.
@@ 7.2.1. Independent Parallel Neighbor Traversal Unit
Both the extremely time-consuming mini-batch sampling and SGB stages involve traversing neighbors of target vertices. As indicated by the analysis in Section 4.2.2, this traversal process can be effectively parallelized across various target vertices. However, Section 5.2.1 and observation 21 underscore that concurrent sampling processes frequently contend for cache resources, leading to notable performance degradation in neighbor sampling tasks. Consequently, it is advisable for researchers to design specialized neighbor traversal units. Each unit should independently cache relevant neighbor information specific to its assigned target vertex, thereby leveraging sampling parallelism while mitigating cache contention. Moreover, optimization structures tailored for graph processing with irregular memory access patterns such as Graphicinado (Ham et al., 2016) and GraphDynS (Yan et al., 2019) could also be contemplated for enhancing neighbor traversal efficiency.
@@ 7.2.3. Multi-lane Architecture Supporting Semantic-graph-level Parallelism
Besides enabling data reuse, multiple semantic graphs offer a form of parallelism unique to traditional GNNs, termed semantic graph parallelism. In HGNN execution, the operations within each semantic graph prior to the SF stage are independent of other graphs, inherently possessing parallelism. Researchers can improve training efficiency by devising multi-channel hardware architectures that facilitate the concurrent execution of various semantic graphs like HiHGNN (Xue et al., 2024a) proposed for HGNN inference acceleration.

### v4

@@ 5.2.2. Bandwidth Contention During Data Loading and Synchronization
Data loading and gradient synchronization are two critical steps in distributed training. 22 The factors limiting the performance of data loading and synchronization in multi-device scenarios are primarily the contention for shared bandwidth. Fig. 14(a) illustrates the trend of PCIe bus communication bandwidth between the host CPU and GPU devices during distributed training as the number of devices increases. Based on the geometric mean across various model and dataset scenarios, an increase in the number of devices from one to two results in an 8.37% reduction in the average PCIe bus bandwidth connecting the host and devices. This reduction further extends to 22.62% when the number of devices increases to four. The decline in PCIe bandwidth suggests that data transfers between the CPU and GPU, particularly the data loading process, encounter increased latency, consequently diminishing the overall performance of distributed training.
@@
Figure 14. Bandwidth contention between devices: (a) Host-Device PCIe bandwidth; (b) Device-Device NVLink bandwidth.
@@
Fig. 14(b) depicts the variation in data transfer bandwidth via NVLink between GPUs during distributed training as the number of devices increases. Based on the geometric mean across various model and dataset configurations, when the number of devices increases from one to two, the average NVLink bandwidth between devices decreases by 26.07%. When the number of devices increases to four, the average NVLink bandwidth further decreases by 46.03%. This reduction can be attributed to the use of the NVIDIA Collective Communications Library for gradient synchronization, which, in this study, employs the All-Reduce algorithm to synchronize gradients across devices. As this algorithm is executed concurrently on multiple devices, the complexity of the communication network increases with the number of devices, leading to intensified contention for available bandwidth among the devices.
@@ 6.1.1. Unique Process of Metapath Instance Generation
Given that HGNNs primarily operate on semantic graphs during execution, it is imperative to construct these semantic graphs from the original HetG, which is unique to HGNNs. Our experimental results for the HAN model on different datasets indicate that the execution time of the SGB stage is, on average, 11.46$\times$ the execution time of one single training epoch. Furthermore, as the length and number of metapaths increase, the execution time of the SGB stage exhibits a marked increase, irrespective of the dataset’s scale. For large-scale HetG datasets, SGB will be a considerably time-consuming stage during the training process.
@@ 6.1.3. Intricate Two-level Aggregation
GNNs perform a single aggregation step for neighboring vertices within a singular type of relation. In contrast, HGNNs aggregate features from neighbors in each semantic graph generated according to corresponding semantics (relations or metapaths), and then fuse intermediate results of each semantic graph for each vertex. As depicted in Fig. 3, the incorporation of a secondary aggregation level, denoted as the SF stage, incurs an approximate 15.38% increase in time overhead per epoch.
@@ 6.2.1. More Intricate Hybrid Execution Pattern
Prior work (Yan et al., 2020) highlights the presence of a hybrid execution pattern in the execution of the typical GCN model. Specifically, the FP stage (Combination stage in work (Yan et al., 2020)) exhibits a more regular pattern and demonstrates compute-bound execution modes, whereas the NA (Aggregation stage) stage involves numerous random accesses, resulting in memory-bound behavior. According to the analysis in Section 4.1.2, compared to GNNs, HGNNs exhibit more pronounced and complex hybrid execution pattern, which stems from their more intricate model structure.
@@ 7. Optimization Guidelines
In this section, leveraging the findings previously outlined, we provide guidance for optimizing HGNN training from both software and hardware perspectives.
7.1. Software Optimizations
@@ 7.1.1. Reasonable Overlapping of Phases
On one hand, a bound-aware kernel fusion method can be proposed to facilitate the overlapping execution of stages with differing bounds. Observations 5 and 6 underscore the presence of intricate hybrid execution patterns during HGNN training. These stages, distinguished by diverse execution bounds, frequently alternate, facilitating overlap execution to harness multiple hardware resources concurrently. This approach enhances overall execution performance by optimizing hardware resource utilization. For example, Graphite (Gong et al., 2022) adopts a similar approach to accelerate GNNs on the CPU.
@@ 7.2.1. Independent Parallel Neighbor Traversal Unit
Both the extremely time-consuming mini-batch sampling and SGB stages involve traversing neighbors of target vertices. As indicated by the analysis in Section 4.2.2, this traversal process can be effectively parallelized across various target vertices. However, Section 5.2.1 and observation 21 underscore that concurrent sampling processes frequently contend for shared resources, leading to notable performance degradation in neighbor sampling tasks. Consequently, it is advisable for researchers to design specialized neighbor traversal units. Each unit should independently cache relevant neighbor information specific to its assigned target vertex, thereby leveraging sampling parallelism while mitigating cache contention. Moreover, optimization structures tailored for graph processing with irregular memory access patterns such as Graphicinado (Ham et al., 2016) and GraphDynS (Yan et al., 2019) could also be contemplated for enhancing neighbor traversal efficiency.
@@ 7.2.2. Unified Reconfigurable Execution Unit
7.2.3. Reuse-distance-aware Cache
Finding 11 reveals that memory dependency is the primary cause of execution stalls in most CUDA kernels during HGNN training, with memory access latency serving as a significant performance bottleneck. Additionally, observation 7 shows that backward propagation incurs more DRAM accesses than forward propagation, while finding 10 reveals a lower cache hit rate during backward passes. This is due to the irregularity of graph data, the broader range of accessed data, and the deeper computational graphs in training, resulting in longer reuse distances and frequent eviction to off-chip memory before reuse. From a holistic perspective, researchers can model the computational graph of HGNN training, and, by considering data size and reuse distance as a function of graph depth, compute a comprehensive replacement weight index. This index can guide the design of specialized cache structures that prioritize retaining data with shorter reuse distances, thereby improving cache hit rates and reducing off-chip memory access.
@@ 7.2.4. Multi-lane Architecture Supporting Semantic-graph-level Parallelism
Besides enabling data reuse, multiple semantic graphs offer a form of parallelism unique to traditional GNNs, termed semantic graph parallelism. During the training process of HGNN models, the operations within each semantic graph prior to the SF stage are independent of other graphs, inherently possessing parallelism. Researchers can improve training efficiency by devising multi-lane hardware architectures that facilitate the concurrent execution of various semantic graphs like HiHGNN (Xue et al., 2024a) proposed for HGNN inference acceleration.